# Community detection for Contexual-LSBM: Theoretical limitation on misclassfication ratio and effecient algorithm

### v1

@@ A.1 Part 1
Leverage Lemma 2 we assert there exist $(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star})$ such that
@@
Recall the definition of the log ratio Eq. 13:
@@
Mimic the approach described in Yun and Proutiere (2016), we could reach inequality $(17)$ in Yun and Proutiere (2016) adopted with our new defined log-ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the necessary condition for $\mathbb{E}[\epsilon(n)]\leq s$:
@@
Then the next part finishes the proof by estiamting the bounds of $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$ respectively.
@@ A.2 Part 2
  • Upper-bound for $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$:
@@
    $\Gamma_{1}$ has been carefully studied in (Yun and Proutiere, 2016, Theorem 1) such that
@@
    Furthermore repeating the procedure in (A.2) leads to
@@
    Then the following part remains to upper bound $\mathbb{E}_{\Psi}[L^{2}]$: Without loss of generaloity we assume $\sigma(i^{\star})=k_{1}$, by definition we have
@@
    Here we use $\mathbb{P}(i^{\star}>\log^{2}(n))\leq n^{-4}$ in the second step, independence of $\{\mathbf{x}\}_{i\in[L]}$ and $\mathbf{A}$ in the third step, equation (33) and assumption 1 in the fourth steps and $10<t<$ 1 in the last step. To proceeds we then focus on $\Gamma_{1}$:
@@
    Here we use the fact that any two nodes $i_{1}$ and $i_{2}$ such that $i_{1}\neq i_{2}\geq i^{\star}$ are muturally independent in the derivation. To upper bound $\Gamma_{11}$ we first have
@@
    We thus finish upper-bounding $\Gamma_{1}$ by analyzing $\Gamma_{11}$:
@@
    Further we end the proof by plugging above inequation back to (A.2):
@@
Our proof then ends here.
@@ Appendix B Proof of Remark 1
Recall the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$, pick any $k_{1}\neq k_{2}\in[K]$ and denote $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ such that
@@
Here $D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})$ and $D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ follow the definition in (3.2). Thus the proof remains to display
@@
We start the proof by studying the cosntrained version of (27): Notice that $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ actually reach the minimum of the following problem:
@@
Thus we write the lagrange dual form of Appendix B such that
@@
for some postive variable $t>0$. Notice that $\Gamma_{1}\left(\mathbf{q}_{A},t\right)$ has been completely studied in claim $4$ in Yun and Proutiere (2016), due to the independence of the $\mathbf{q}_{A},q_{X}$, the remaining part then focues on solving $\Gamma_{2}(q_{X},t)$.
Since $q_{X}$ represent a probability density function that lies in the function space $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Denote $\frac{\delta F(p)}{\delta p}(\mathbf{x})$ as the functional derivative of $F(p)$ to $p$, $\forall p\in\mathcal{F}$, by definition we have
@@
Here we use $\frac{\delta\operatorname{KL}(q,p)}{\delta q}(\mathbf{x})=\log(q(\mathbf{x}))+1-\log(p(\mathbf{x}))$ in the derivation. Denote $q_{X}^{\star}(x):=\operatorname*{argmin}_{q_{X}\in\mathcal{F}}\Gamma_{2}(q_{X},t)$ and let (B) equal to zero we have:
@@
due to the convexity of $\Gamma_{2}(q_{X},t)$. Notice that $p_{X}(k_{1})$ and $p_{X}(k_{2})$ represent the pdf for $\mathcal{N}(\boldsymbol{\mu}_{k_{1}},\mathbf{I}_{d})$ and $\mathcal{N}(\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ respectively, consequently we obtain
@@
after some calculation. Here $C$ is some constant that is irelevent with $\mathbf{x}$. We assert $C=(2\pi)^{-d/2}$ by utilzing $q_{X}$ is an probability density function.
Plug the expression of $q_{X}^{\star}$ back to $\Gamma_{2}(q_{X},t)$ defined in (B) gives:
@@
after some algebra. Here we use Lemma 3 and the fact that $q_{X}^{\star}$ is the PDF for $\mathcal{N}((1-t)\boldsymbol{\mu}_{k_{1}}+t\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ in the derivation. We thus finish the proof by combining the expression of $\Gamma_{1}(\mathbf{q}_{A},t)$(see (Yun and Proutiere, 2016, Claim $4$)) and $\Gamma_{2}(q_{X}^{\star},t)$ together.
@@ Appendix C Proof of Lemma 1
with probability at least $1-cn^{-c^{\prime}}$. Here we use $\mathbf{S}^{(k)}$ is the best approximation of $\mathbf{S}$ in the third step, (Coja-Oghlan, 2010, Lemma 8.5) and (Löffler et al., 2021, Lemma B.1) in the last step. Notice that $\operatorname{rank}~{}(\mathbf{S}^{(k)}-\mathbf{M})\leq\operatorname{rank}(\mathbf{S}^{(k)})+\operatorname{rank}(\mathbf{M})\leq 2K$ we further have
@@
where $\boldsymbol{\Sigma}:=\operatorname{diag}([n_{1},n_{2},\cdots,n_{K}])=n\left(\operatorname{diag}(\boldsymbol{\alpha})+o(1)\mathbf{I}_{K}\right)$. Here we use the orthonormality for the columns of $\mathbf{Z}\boldsymbol{\Sigma}^{-1}$ in the last step. Thus we further obtain
@@ Appendix D Technical lemmas
Consider two Gaussian probability distributions with unit variance, having means $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and probability density functions (pdfs) denoted as $p_{1}$ and $p_{2}$, respectively. It follows that one has
@@
Recall the definition of $\bar{\sigma}(\cdot)$ in (1) we have $\bar{\sigma}(\pi(i))=\hat{\sigma}(i)$ for some permutation $\pi(\cdot)$.

### v3

@@ A.1 Part 1
Leverage Lemma 2 we assert there exist $(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star})$ such that:
@@
Recall the definition of the log ratio Eq. 13 we obtain:
@@
Mimicking the approach described in Yun and Proutiere (2016), we can derive inequality $(17)$ in Yun and Proutiere (2016) adapted with our new defined log-ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the necessary condition for $\mathbb{E}[\epsilon(n)]\leq s$:
@@
Then the next part finishes the proof by estimating the bounds of $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$ respectively.
@@ A.2 Part 2
  • Upper bound for $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$:
@@
    $\Gamma_{1}$ has been carefully studied in (Yun and Proutiere, 2016, Theorem 1) such that:
@@
    Furthermore, repeating the procedure in (A.2) leads to
@@
    Thus, the following part remains to upper bound $\mathbb{E}_{\Psi}[L^{2}]$: Without loss of generality we assume $\sigma(i^{\star})=k_{1}$, by definition we have
@@
    Here, we use $\mathbb{P}(i^{\star}>\log^{2}(n))\leq n^{-4}$ in the second step, the independence of $\{\mathbf{x}\}_{i\in[L]}$ and $\mathbf{A}$ in the third step, equation (33) and 1 in the fourth step and $10<t<$ 1 in the last step. To proceeds we then focus on $\Gamma_{1}$:
@@
    Here we use the fact that any two nodes $i_{1}$ and $i_{2}$ satisfying $i_{1}\neq i_{2}\geq i^{\star}$ are mutually independent in the derivation. To upper bound $\Gamma_{11}$, we first have
@@
    We thus complete the upper bound of $\Gamma_{1}$ by analyzing $\Gamma_{11}$:
@@
    Further we end the proof of upper bounding by plugging above inequalities back to (A.2):
@@
Our proof concludes here.
@@ Appendix B Proof of Remark 1
Recall the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$. For any $k_{1}\neq k_{2}\in[K]$ define $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ such that:
@@
Here, $D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})$ and $D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ follow the definition in (3.2). Thus the proof reduces to showing:
@@
We begin the proof by studying the constrained version of (27): Note that $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ actually reaches the minimum of the following problem:
@@
Thus, we express the lagrange dual form of Appendix B such that:
@@
for some postive variable $t>0$. Notice that $\Gamma_{1}\left(\mathbf{q}_{A},t\right)$ has been completely studied in Claim $4$ in Yun and Proutiere (2016), due to the independence of the $\mathbf{q}_{A},q_{X}$, the remaining part then focuses on solving $\Gamma_{2}(q_{X},t)$.
Since $q_{X}$ represents a PDF that lies in the function space $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. We denote $\frac{\delta F(p)}{\delta p}(\mathbf{x})$ as the functional derivative of $F(p)$ with respect to $p$, $\forall p\in\mathcal{F}$. By definition, we have
@@
Here we use $\frac{\delta\operatorname{KL}(q,p)}{\delta q}(\mathbf{x})=\log(q(\mathbf{x}))+1-\log(p(\mathbf{x}))$ in the derivation. Let $q_{X}^{\star}(x):=\operatorname*{argmin}_{q_{X}\in\mathcal{F}}\Gamma_{2}(q_{X},t)$. Setting (B) equal to zero, we have:
@@
due to the convexity of $\Gamma_{2}(q_{X},t)$. Note that $p_{X}(k_{1})$ and $p_{X}(k_{2})$ represent the PDF for $\mathcal{N}(\boldsymbol{\mu}_{k_{1}},\mathbf{I}_{d})$ and $\mathcal{N}(\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ respectively. Consequently, after some calculation, we obtain
@@
after some calculation. Here $C$ is a constant that is irelevent with $\mathbf{x}$. We assert $C=(2\pi)^{-d/2}$ by utilizing $q_{X}$ is an probability density function.
Substituting the expression of $q_{X}^{\star}$ into $\Gamma_{2}(q_{X},t)$ defined in (B) yields:
@@
after some algebra. Here, we use Lemma 3 and the fact that $q_{X}^{\star}$ is the PDF for $\mathcal{N}((1-t)\boldsymbol{\mu}_{k_{1}}+t\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ in the derivation. We thus finish the proof by combining the expressions of $\Gamma_{1}(\mathbf{q}_{A},t)$(see (Yun and Proutiere, 2016, Claim $4$)) and $\Gamma_{2}(q_{X}^{\star},t)$ together.
@@ Appendix C Proof of Lemma 1
with probability at least $1-cn^{-c^{\prime}}$. Here we use tha fact that $\mathbf{S}^{(K)}$ is the best approximation of $\mathbf{S}$ in the third step, (Coja-Oghlan, 2010, Lemma 8.5) and (Löffler et al., 2021, Lemma B.1) in the last step. Notice that $\operatorname{rank}~{}(\mathbf{S}^{(K)}-\mathbf{M})\leq\operatorname{rank}(\mathbf{S}^{(K)})+\operatorname{rank}(\mathbf{M})\leq 2K$ we further have
@@
where $\boldsymbol{\Sigma}:=\operatorname{diag}([n_{1},n_{2},\cdots,n_{K}])=n\left(\operatorname{diag}(\boldsymbol{\alpha})+\mathbf{\Delta})\mathbf{I}_{K}\right)$ for some $\mathbf{\Delta}$ i.e. $\left\|\mathbf{\Delta}\right\|_{op}=o(1)$. Here we use the orthonormality for the columns of $\mathbf{Z}\boldsymbol{\Sigma}^{-1}$ in the last step. Thus we further obtain
@@ Appendix D Technical lemmas
Consider two Gaussian probability distributions with unit variance, having means $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and PDFs denoted as $p_{1}$ and $p_{2}$, respectively. It follows that one has
@@
Recall the definition of $\bar{\sigma}(\cdot)$ in (1) we have $\pi(\bar{\sigma}(i))=\hat{\sigma}(i)$ for some permutation $\pi(\cdot)$.